For each of the following equations you should locate all equilibrium points and investigate their stability properties for ddi = (x – 1) (u + (x + 1)2)
Answer #1

solution 8 Given that Let f(x) =(x-1) (4+ (+) Let y = x-1. This geven equation becomes Ý =ý (Mtoty? Let f (y) = y put i+y3 9=.y=0 (X=1) ys + Fi-ue :(4x = 10-i-ue) tco) Eltico, so yoo is stable fi I T ) =M+1+36-1-4)=-2-2470 1 so, Isoleu as unstable Wso at M = utal, there is Pitchfork, bituscation. +* HA - stable HH unstabde. 9HHHHHHHHH Eelari bifoscation diagram O-